# Optimally truncated WKB approximation for the 1D stationary Schrödinger equation in the highly oscillatory regime

### v1

@@ 5.1 Example 1: Airy equation
Next, we investigate numerically the behavior of the WKB approximation error $\lVert\varphi-\widetilde{\varphi}_{N}^{WKB}\rVert_{L^{\infty}(I)}$ as a function of $\varepsilon$. We may compare the results with the error “estimate” (77). As a first test, we set $M=8$ to compute $\widetilde{\varphi}_{N}^{WKB}$. This results in an approximation error for $S_{n}$ of $e_{n}\approx 10^{-8}$, $n=0,\dots,4$. On the left of Figure 3 we plot for $N=0,\dots,4$ the error as a function of $\varepsilon$: For $N=2,3,4$ and small values of $\varepsilon$, the $\mathcal{O}(\varepsilon^{-1})e_{0}$-term is dominant. In contrast, for $N=0$ and $N=1$ this error term is not visible for the given range of $\varepsilon$-values such that the $\mathcal{O}(\varepsilon^{N})$-term is dominant. As a second test, we set again $M=8$, but now use in $\widetilde{\varphi}_{N}^{WKB}$ the exactly computed function $S_{0}$. The $\mathcal{O}(\varepsilon^{-1})e_{0}$-term from (77) is thus eliminated. On the right of Figure 3 we show again the error $\lVert\varphi-\widetilde{\varphi}_{N}^{WKB}\rVert_{L^{\infty}(I)}$ as a function of $\varepsilon$: For $N=2,3,4$ and small $\varepsilon$-values, the $\mathcal{O}(\varepsilon^{0})e_{1}$-term, which is the next term in the sum in (77), now dominates. Indeed, the error curves show an almost constant value of approximately $2\cdot 10^{-9}$ for small values of $\varepsilon$. For larger $\varepsilon$, the error curves behave like $\mathcal{O}(\varepsilon^{N})$. As a third test, we set $M=25$ and approximate again all functions $S_{n}$, $n=0,\dots,4$ (as in the first test). The corresponding approximation errors of $S_{n}$ are $e_{n}\approx 10^{-23}$, $n=0,\dots,4$. On the left of Figure 4 we present the resulting WKB approximation errors. We observe that, on this scale, all $\mathcal{O}(\varepsilon^{n-1})e_{n}$-terms in the sum of (77) are essentially eliminated, since all the shown error curves behave like $\mathcal{O}(\varepsilon^{N})$. Overall, we observe very good agreement between the numerical results of each of the three tests and the statements from Theorem 3.9 and Remark 3.10.
@@
Figure 3: $L^{\infty}(I)$-norm of the error of the WKB approximation as a function of $\varepsilon$, for the IVP (93) and several choices of $N$. Left: $M=8$. Right: $M=8$; using the exactly computed function $S_{0}$.
@@
Figure 4: Left: $L^{\infty}(I)$-norm of the error of the WKB approximation as a function of $\varepsilon$, for the IVP (93) and several choices of $N$. Here, we set $M=25$. Right: $L^{\infty}(I)$-norm of the error of the WKB approximation as a function of $N$, for the IVP (93) and several choices of $\varepsilon$. The dash-dotted lines correspond to the error estimate according to Theorem 3.7 and the solid lines correspond to the actual error of the WKB approximation.
@@
Next we investigate the error of the (perturbed) WKB approximation as a function of the truncation order $N$. For this, we set again $M=25$, yielding approximation errors of $S_{n}$ as $e_{n}\approx 10^{-23}$ for $n=0,1,\dots$. We may therefore neglect the errors caused by approximating the functions $S_{n}$. On the right of Figure 4 we plot the actual error $\lVert\varphi-\widetilde{\varphi}_{N}^{WKB}\rVert_{L^{\infty}(I)}$ and its error estimate (3.7) while again using $K_{2}^{fitted}=10/37$, both as functions of $N$, for several $\varepsilon$-values. We observe that, even when using the fitted constant $K_{2}^{fitted}$, the “optimal” truncation order $\widehat{N}_{opt}$, as predicted by the estimate (3.7), is smaller than $N_{opt}$ (determined as the $\operatorname{argmin}$ of the actual error curve). For instance, we have $\widehat{N}_{opt}(2^{-4})\approx 14<22\approx N_{opt}(2^{-4})$ and $\widehat{N}_{opt}(2^{-5})\approx 29<44\approx N_{opt}(2^{-5})$, respectively. In Figure 5 we plot on the left the optimal truncation order $N_{opt}(\varepsilon)$ as a function of $\varepsilon$ as well as its predictions $\widehat{N}_{opt}(\varepsilon)$, $N_{heu}(\varepsilon)$, and $\widehat{N}_{heu}(\varepsilon)$. The plot suggests that $N_{opt}$, $\widehat{N}_{opt}$, and $N_{heu}$ are proportional to $\varepsilon^{-1}$, as $\varepsilon\to 0$ (for $\widehat{N}_{heu}$ this is already evident from (84)). Further, for $\varepsilon=2^{-1},2^{-3},2^{-4},2^{-5}$ we observe that $N_{opt}=N_{heu}$. On the right of Figure 5 we plot the corresponding optimal error which is achieved by using $N_{opt}$ as well as error estimate (3.7) when using $N=\widehat{N}_{opt}$, both as a function of $\varepsilon$. As indicated by the dashed line, the optimal error decays like $\mathcal{O}(\varepsilon^{-2}\exp(-r/\varepsilon))$, with $r\approx 13/10$ being a fitted value, in good agreement with Remark 4.2.
@@
Figure 5: Left: The optimal truncation order $N_{opt}$ as well as the predicted “optimal” orders $\widehat{N}_{opt}$, $N_{heu}$, and $\widehat{N}_{heu}$ as functions of $\varepsilon$. The dashed line is proportional to $1/\varepsilon$. Right: The optimal error achieved by using $N_{opt}$ as well as error estimate (3.7) when using $N=\widehat{N}_{opt}$, both as functions of $\varepsilon$. The dashed line is proportional to $\varepsilon^{-2}\exp(-\frac{13}{10\varepsilon})$.
@@ 5.2 Example 2
Here, $J_{\nu}$ and $Y_{\nu}$ denote the Bessel functions of first and second kind of order $\nu$, respectively (e.g., see [21, Chap. 10]).
@@
Figure 6: Left: Exact solution (97) of IVP (96) for the choice $\varepsilon=10^{-2}$. Right: $L^{\infty}(I)$-norm of $S_{n}^{\prime}$ and $\widetilde{S}_{n}$ as functions of $n$, for the example $a(x)=\exp(5x)$ on the interval $I=[0,1]$.
Further, we plot on the right of Figure 6 the $L^{\infty}(I)$-norms of the derivatives $S_{n}^{\prime}$ and the approximations $\widetilde{S}_{n}$ when using $M=30$. As indicated by the dashed line, the smallest (fitted) constant $K_{2}$ such that estimate (39) holds is $K_{2}^{fitted}\approx 9/20$ (here we also replaced $\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}$ in (39) by the fitted value $0.6$). In Figure 7 we present the WKB approximation error $\lVert\varphi-\widetilde{\varphi}_{N}^{WKB}\rVert_{L^{\infty}(I)}$ as a function of $\varepsilon$ and may again compare the results with the error “estimate” (77). We observe that, on this scale, all $\mathcal{O}(\varepsilon^{n-1})e_{n}$-terms are essentially eliminated, since all the shown error curves behave like $\mathcal{O}(\varepsilon^{N})$. Overall, we observe very good agreement with the statements from Theorem 3.9 and Remark 3.10. Finally, we plot in Figure 8 on the left the optimal truncation order $N_{opt}$ as well as its predictions $\widehat{N}_{opt}$, $N_{heu}$, and $\widehat{N}_{heu}$, as functions of $\varepsilon$. We find that $N_{opt}$ is proportional to $\varepsilon^{-1}$, as $\varepsilon\to 0$. On the right of Figure 8 we present the corresponding optimal error as well as error estimate (3.7) when using $N=\widehat{N}_{opt}$, both as a function of $\varepsilon$. As the dashed line indicates, the error decays like $\mathcal{O}(\varepsilon^{-2}\exp(-r/\varepsilon))$, with $r\approx 3/4$ being a fitted value. This is in good agreement with Remark 4.2.
Figure 7: $L^{\infty}(I)$-norm of the error of the WKB approximation as a function of $\varepsilon$, for the IVP (96) and several choices of $N$. Here, we set $M=30$.
Figure 8: Left: The optimal truncation order $N_{opt}$ as well as the predicted “optimal” orders $\widehat{N}_{opt}$, $N_{heu}$, and $\widehat{N}_{heu}$ as functions of $\varepsilon$. The dashed line is proportional to $1/\varepsilon$. Right: The optimal error achieved by using $N_{opt}$ as well as error estimate (3.7) when using $N=\widehat{N}_{opt}$, both as functions of $\varepsilon$. The dashed line is proportional to $\varepsilon^{-2}\exp(-\frac{3}{4\varepsilon})$.
@@ Appendix A Convergent WKB series
In this appendix we provide examples where the asymptotic series (6) is convergent in $L^{\infty}(I)$. In practice, the norms $\lVert S_{n}^{\prime}\rVert_{L^{\infty}(I)}$ (and $\lVert S_{n}\rVert_{L^{\infty}(I)}$) often decrease up to a certain number of $n$ before they start to increase rapidly, e.g., see the right plot of Figure 2. However, there are examples where one can easily verify that this is not the case. For instance, consider the simplest case in which $a\equiv a_{0}$ is constant. By (9) this is equivalent to $S_{1}^{\prime}\equiv 0$, which by (10) then implies $S_{n}^{\prime}\equiv 0$ for every $n\geq 1$. Similarly, one easily verifies that $S_{2}^{\prime}\equiv 0$ is equivalent to $a$ having the form $a(x)=(C_{1}+C_{2}x)^{-4}$ for some constants $C_{1}$ and $C_{2}$, see also [5, Problem 10.2]. It then holds $S_{n}^{\prime}\equiv 0$ for every $n\geq 2$. Thus, in both of the just mentioned cases, the asymptotic series (6) terminates automatically and is therefore convergent. The corresponding WKB approximation (13) with $N\geq 0$ (respectively $N\geq 1$) is then the exact solution to IVP (1). Indeed, revisiting (48), it is clear that the r.h.s. in (3) then vanishes, i.e. $\lVert\varphi-\varphi_{N}^{WKB}\rVert_{L^{\infty}(I)}=0$.
@@
Thus, the constants $C_{i}$, $i=1,2,3$, determine whether the function $\lVert S_{2n}\rVert_{L^{\infty}(I)}$ is exponentially growing or decaying, as $n\to\infty$. Note that Proposition A.1 also implies that $S_{2n+1}\equiv 0$ for $n\geq 1$. A short calculation then shows that (109) implies that the corresponding WKB series $\exp(\sum_{n=0}^{\infty}\varepsilon^{n-1}S_{n}(x))$ is (geometrically) convergent for any $x\in I$, if $\varepsilon\leq|C_{1}C_{3}-C_{2}^{2}/4|^{-1/2}$.

### v2

@@ 5.1 Example 1: Airy equation
Next, we investigate numerically the behavior of the WKB approximation error $\lVert\varphi-\widetilde{\varphi}_{N}^{WKB}\rVert_{L^{\infty}(I)}$ as a function of $\varepsilon$. We may compare the results with the error “estimate” (62). As a first test, we set $M=8$ to compute $\widetilde{\varphi}_{N}^{WKB}$. This results in an approximation error for $S_{n}$ of $e_{n}\approx 10^{-8}$, $n=0,\dots,4$. On the left of Figure 4 we plot for $N=0,\dots,4$ the error as a function of $\varepsilon$: For $N=2,3,4$ and small values of $\varepsilon$, the $\mathcal{O}(\varepsilon^{-1})e_{0}$-term is dominant. In contrast, for $N=0$ and $N=1$ this error term is not visible for the given range of $\varepsilon$-values so that the $\mathcal{O}(\varepsilon^{N})$-term is dominant. As a second test, we set again $M=8$, but now use in $\widetilde{\varphi}_{N}^{WKB}$ the exactly computed function $S_{0}$. The $\mathcal{O}(\varepsilon^{-1})e_{0}$-term from (62) is thus eliminated. On the right of Figure 4 we show again the error $\lVert\varphi-\widetilde{\varphi}_{N}^{WKB}\rVert_{L^{\infty}(I)}$ as a function of $\varepsilon$: For $N=2,3,4$ and small $\varepsilon$-values, the $\mathcal{O}(\varepsilon^{0})e_{1}$-term, which is the next term in the sum in (62), now dominates. Indeed, the error curves show an almost constant value of approximately $2\cdot 10^{-9}$ for small values of $\varepsilon$. For larger $\varepsilon$, the error curves behave like $\mathcal{O}(\varepsilon^{N})$. As a third test, we set $M=25$ and approximate again all functions $S_{n}$, $n=0,\dots,4$ (as in the first test). The corresponding approximation errors of $S_{n}$ are $e_{n}\approx 10^{-23}$, $n=0,\dots,4$. On the left of Figure 5 we present the resulting WKB approximation errors. We observe that, on this scale, all $\mathcal{O}(\varepsilon^{n-1})e_{n}$-terms in the sum of (62) are essentially eliminated, since all the shown error curves behave like $\mathcal{O}(\varepsilon^{N})$. Overall, we observe very good agreement between the numerical results of each of the three tests and the statements from Theorem 3.9 and Remark 3.10.
@@
Figure 4: $L^{\infty}(I)$-norm of the error of the WKB approximation as a function of $\varepsilon$, for the IVP (78) and several choices of $N$. Left: $M=8$. Right: $M=8$; using the exact function $S_{0}$.
@@
Figure 5: Left: $L^{\infty}(I)$-norm of the error of the WKB approximation as a function of $\varepsilon$ for the IVP (78) and several choices of $N$. Here, we set $M=25$. Right: $L^{\infty}(I)$-norm of the error of the WKB approximation as a function of $N$ for the IVP (78) and several choices of $\varepsilon$. The dash-dotted lines correspond to the error estimate according to Theorem 3.7 and the solid lines correspond to the actual error of the WKB approximation. (In both plots the curves have the same top-down ordering as the legend.)
@@
Next we investigate the error of the (perturbed) WKB approximation as a function of the truncation order $N$. For this, we set again $M=25$, yielding approximation errors of $S_{n}$ as $e_{n}\approx 10^{-23}$ for $n=0,1,\dots$. We may therefore neglect the errors caused by approximating the functions $S_{n}$. On the right of Figure 5 we plot the actual error $\lVert\varphi-\widetilde{\varphi}_{N}^{WKB}\rVert_{L^{\infty}(I)}$ and its error estimate (3.7) while again using $K_{2}^{fitted}=10/37$, both as functions of $N$, for several $\varepsilon$-values. We observe that, even when using the fitted constant $K_{2}^{fitted}$, the “optimal” truncation order $\widehat{N}_{opt}$, as predicted by the estimate (3.7), is smaller than $N_{opt}$ (determined as the $\operatorname{argmin}$ of the actual error curve). For instance, we have $\widehat{N}_{opt}(2^{-4})\approx 14<22\approx N_{opt}(2^{-4})$ and $\widehat{N}_{opt}(2^{-5})\approx 29<44\approx N_{opt}(2^{-5})$, respectively. This is not a paradox, but it is implied in this example by the strong over-estimation (3.7) of the error for large $N$.
In Figure 6 we plot on the left the optimal truncation order $N_{opt}(\varepsilon)$ as a function of $\varepsilon$ as well as its predictions $\widehat{N}_{opt}(\varepsilon)$, $N_{heu}(\varepsilon)$, and $\widehat{N}_{heu}(\varepsilon)$. The plot suggests that $N_{opt}$, $\widehat{N}_{opt}$, and $N_{heu}$ are all proportional to $\varepsilon^{-1}$, as $\varepsilon\to 0$ (for $\widehat{N}_{heu}$ this is already evident from (69)). Further, for $\varepsilon=2^{-1},2^{-3},2^{-4},2^{-5}$ we observe that $N_{opt}=N_{heu}$. On the right of Figure 6 we plot the corresponding optimal error which is achieved by using $N_{opt}$ as well as error estimate (3.7) when using $N=\widehat{N}_{opt}$, both as a function of $1/\varepsilon$. As indicated by the dashed line, the optimal error decays like $\mathcal{O}(\exp(-r/\varepsilon))$, with $r\approx 1.36$ being a fitted value, in good agreement with Remark 4.2.
@@
Figure 6: Left: The optimal truncation order $N_{opt}$ as well as the predicted “optimal” orders $\widehat{N}_{opt}$, $N_{heu}$, and $\widehat{N}_{heu}$ as functions of $\varepsilon$. The dashed line is proportional to $1/\varepsilon$. Right: The optimal error achieved by using $N_{opt}$ as well as error estimate (3.7) when using $N=\widehat{N}_{opt}$, both as functions of $1/\varepsilon$. The dashed line is proportional to $\exp(-\frac{r}{\varepsilon})$ with $r\approx 1.36$.
@@ 5.2 Example 2
Here, $J_{\nu}$ and $Y_{\nu}$ denote the Bessel functions of first and second kind of order $\nu$, respectively (e.g., see [24, Chap. 10]).
@@
Figure 7: Left: Exact solution (82) of IVP (81) for the choice $\varepsilon=10^{-2}$. Right: $L^{\infty}(I)$-norm of $S_{n}^{\prime}$ and $\widetilde{S}_{n}$ as functions of $n$ for the example $a(x)=\exp(5x)$ on the interval $I=[0,1]$.
Further, we plot on the right of Figure 7 the $L^{\infty}(I)$-norms of the derivatives $S_{n}^{\prime}$ and the approximations $\widetilde{S}_{n}$ when using $M=30$. As indicated by the dashed line, the smallest (fitted) constant $K_{2}$ such that estimate (39) holds is $K_{2}^{fitted}\approx 9/20$ (here we also replaced $\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}$ in (39) by the fitted value $0.6$). In Figure 8 we present the WKB approximation error $\lVert\varphi-\widetilde{\varphi}_{N}^{WKB}\rVert_{L^{\infty}(I)}$ as a function of $\varepsilon$ and may again compare the results with the error “estimate” (62). We observe that, on this scale, all $\mathcal{O}(\varepsilon^{n-1})e_{n}$-terms are essentially eliminated, since all the shown error curves behave like $\mathcal{O}(\varepsilon^{N})$. Overall, we observe very good agreement with the statements from Theorem 3.9 and Remark 3.10. Finally, we plot in Figure 9 on the left the optimal truncation order $N_{opt}$ as well as its predictions $\widehat{N}_{opt}$, $N_{heu}$, and $\widehat{N}_{heu}$, as functions of $\varepsilon$. We find that $N_{opt}$ is proportional to $\varepsilon^{-1}$, as $\varepsilon\to 0$. On the right of Figure 9 we present the corresponding optimal error as well as error estimate (3.7) when using $N=\widehat{N}_{opt}$, both as a function of $1/\varepsilon$. As the dashed line indicates, the error decays like $\mathcal{O}(\exp(-r/\varepsilon))$, with $r\approx 0.81$ being a fitted value. This is in good agreement with Remark 4.2.
Figure 8: $L^{\infty}(I)$-norm of the error of the WKB approximation as a function of $\varepsilon$ for the IVP (81) and several choices of $N$. Here, we set $M=30$.
Figure 9: Left: The optimal truncation order $N_{opt}$ as well as the predicted “optimal” orders $\widehat{N}_{opt}$, $N_{heu}$, and $\widehat{N}_{heu}$ as functions of $\varepsilon$. The dashed line is proportional to $1/\varepsilon$. Right: The optimal error achieved by using $N_{opt}$ as well as error estimate (3.7) when using $N=\widehat{N}_{opt}$, both as functions of $1/\varepsilon$. The dashed line is proportional to $\exp(-\frac{r}{\varepsilon})$ with $r\approx 0.81$.
@@ Appendix A Convergent WKB series
In this appendix we provide examples where the asymptotic series (6) is convergent in $L^{\infty}(I)$. In practice, the norms $\lVert S_{n}^{\prime}\rVert_{L^{\infty}(I)}$ (and $\lVert S_{n}\rVert_{L^{\infty}(I)}$) often decrease up to a certain number of $n$ before they start to increase rapidly, e.g., see the right plot of Figure 3. However, there are examples where one can easily verify that this is not the case. For instance, consider the simplest case in which $a\equiv a_{0}$ is constant. By (9) this is equivalent to $S_{1}^{\prime}\equiv 0$, which by (10) then implies $S_{n}^{\prime}\equiv 0$ for every $n\geq 1$. Similarly, one easily verifies that $S_{2}^{\prime}\equiv 0$ is equivalent to $a$ having the form $a(x)=(C_{1}+C_{2}x)^{-4}$ for some constants $C_{1}$ and $C_{2}$, see also [6, Problem 10.2]. It then holds $S_{n}^{\prime}\equiv 0$ for every $n\geq 2$. Thus, in both of the just mentioned cases, the asymptotic series (6) terminates automatically and is therefore convergent. The corresponding WKB approximation (13) with $N\geq 0$ (respectively $N\geq 1$) is then the exact solution to IVP (1). Indeed, revisiting (43), it is clear that the r.h.s. in (3) then vanishes, i.e. $\lVert\varphi-\varphi_{N}^{WKB}\rVert_{L^{\infty}(I)}=0$.
@@
Thus, the constants $C_{i}$, $i=1,2,3$, determine whether the function $\lVert S_{2n}\rVert_{L^{\infty}(I)}$ is exponentially growing or decaying as $n\to\infty$. Note that Proposition A.1 also implies that $S_{2n+1}\equiv 0$ for $n\geq 1$. A short calculation then shows that (94) implies that the corresponding WKB series $\exp(\sum_{n=0}^{\infty}\varepsilon^{n-1}S_{n}(x))$ is (geometrically) convergent for any $x\in I$, if $\varepsilon\leq|C_{1}C_{3}-C_{2}^{2}/4|^{-1/2}$.